# Learning and composing of classical music using restricted Boltzmann machines

### v1

@@ 3.4 Analysis of the Hidden Layer Using t-SNE
Among the 58 compositions used for training, we selected two pieces, BWV857 and BWV868, and prepared transposed versions shifted by a semitone and a whole tone, respectively. For each version, piano roll images were generated, and multiple vectors corresponding to two-measure segments were extracted. These vectors were then provided as inputs to the visible layer of the trained RBM, and the corresponding hidden unit activations were sampled.
Figure 9 shows the result of dimensionality reduction of the hidden unit activations for each input, performed using t-SNE [33]. When transposed by a semitone, the hidden representations corresponding to the original and transposed versions of the same piece were distributed relatively far apart (Fig. 9(a)). If the trained RBM recognized the original and semitone-transposed versions of a musical piece as the same composition, similar to human perception, then the corresponding high-dimensional hidden state vectors would be expected to be projected to nearby locations in the dimensionality-reduced space. However, the fact that the hidden states change significantly upon transposition indicates that the RBM perceives the transposed piece as an entirely different input.
To further examine the criteria by which the trained RBM evaluates musical similarity, we also analyzed the hidden states corresponding to musical data transposed up by a whole tone. The results are shown in Fig. 9(b). In contrast to the case of semitone transposition, the hidden states corresponding to the same original piece appear to be distributed in relatively close proximity. These results suggest that the trained RBM employs a notion of musical similarity that differs from that of human perception.
If the trained RBM recognizes the original music data and the music data transposed up by a semitone as the same, in a manner similar to humans, it would be expected that the pre- and post-transposition data would be plotted close to each other on the graph obtained by dimension reduction of the high-dimensional vectors in the hidden layer. However, since the pre- and post-transposition data were separated on the graph when the music was transposed up by a semitone, it is hypothesized that the trained RBM uses a different metric than humans for recognizing the similarity between musical pieces.
@@
To understand how the similarity between musical keys affects the internal representation of the RBM, we analyzed the number of shared scale tones between the original and transposed keys. For example, in the case of typical seven-note scales such as major and natural minor scales, transposing a piece up by a semitone results in only two common scale tones with the original key. In contrast, transposing it up by a whole tone yields five shared tones. Therefore, whole-tone transposition retains more scale tones in common with the original key than semitone transposition. This suggests a tendency for the hidden state vectors obtained from the RBM to be located closer to each other in the t-SNE projection when the transposed and original pieces share more scale tones.
@@
These results imply that the RBM is more sensitive to the absolute pitch content rather than the relative pitch relationships of a musical input when assessing similarity. In training the RBM, we augmented the dataset by including transpositions of the original pieces up to a major sixth and down to a perfect fourth. This procedure was analogous to data augmentation in image processing, intended to encourage the model to acquire translation invariance. However, the results demonstrate that the trained RBM remains highly sensitive to pitch translations and fails to achieve invariance to transposition. The vulnerability of RBMs to translations of input data has long been pointed out [34], and a similar result was observed in our study using RBMs trained on piano roll data.
@@
4 Summary and Discussion
We demonstrated that music composition is feasible even with a structurally simple model such as the RBM. By adopting the piano roll format as the data representation, we enabled the RBM to learn musical features using methods analogous to those employed for image modeling. The trained RBM was able to successfully reconstruct piano rolls, even for musical pieces that were not included in the training data. We also confirmed that the RBM fails to reconstruct images that are not piano rolls, and correspondingly assigns high energy to such inputs. Although the training data consisted of piano rolls corresponding to two measures, we developed an algorithm that allowed the model to generate musical compositions longer than two measures.
The simplicity of the RBM architecture enables us to analyze how the trained model interprets musical data more easily than with more complex models. To investigate how the RBM internally represents musical information, we analyzed the states of the hidden layer in response to various input data. The results suggest that transposition involves substantial changes in the hidden states, indicating that the RBM tends to assess musical similarity based more on the overlap of absolute pitch positions rather than melodic structure. It has been pointed out that RBMs and Deep Belief Networks (DBNs) lack invariance to translations in the input space. This limitation may explain the observed sensitivity to transpositions. In contrast, Convolutional Deep Belief Models, which incorporate local receptive fields and weight sharing, have the potential to recognize transposed versions of the same music as similar, due to their improved translational invariance [34].
As a future direction, we plan to investigate whether training the RBM on compositions by composers other than J. S. Bach leads to the generation of piano roll images that exhibit different musical characteristics. This would allow us to examine whether the RBM is capable of extracting and reproducing the stylistic features specific to individual composers during the learning process. It has also been reported that the task learned by an RBM is related to the singular values of its weight matrix [35]. Therefore, it would be an intriguing direction for future work to investigate how factors such as the composer or musical genre affect the singular value spectrum of the RBM’s weight matrix.
There have been reports suggesting that each unit in the hidden layer of an RBM encodes a prototypical pattern, such as strokes in figures, in the visible layer [10]. However, such patterns are often not clearly interpretable in standard RBMs. In contrast, Classification RBMs tend to exhibit more distinguishable and meaningful hidden-unit activations [36]. Therefore, using Boltzmann machines with modified architectures may allow us to identify prototypical melodies or chords learned by the model. This remains an important direction for future research.
{acknowledgment}
This research was supported by JSPS KAKENHI, Grant No. JP21K11923. The computation was partly carried out using the facilities of the Supercomputer Center, Institute for Solid State Physics (ISSP), University of Tokyo
Appendix A Training Time of RBM on MNIST
We developed a lightweight Python library that allows easy experimentation with RBMs. The library is available online [30]. It performs computations using NumPy when running on a CPU and switches to CuPy to accelerate processing when a GPGPU is available. To evaluate its performance, we measured the training time of RBMs on the MNIST dataset using System C, a supercomputer at the Institute for Solid State Physics, The University of Tokyo. We measured the computation time both with and without using a GPGPU and compared the results. The CPU used was an AMD EPYC 7763 (2.45 GHz, 64 cores) with 256 GB of memory. For the GPU, we used four NVIDIA A100s, each with 40 GB of memory, totaling 160 GB of GPU memory. The results of plotting the computation time required for training RBMs with different numbers of hidden units are shown in Fig. 10. These results indicate that using a GPGPU significantly reduces computation time.
Figure 10 presents the computation time required to train RBMs with various numbers of hidden units. The results clearly demonstrate that GPGPU-based training significantly reduces the overall computation time compared to CPU-only execution.

### v3

@@ Appendix A Analysis of the Hidden Layer Using t-SNE
Among the 58 compositions used for training, two pieces (BWV857 and BWV868) were selected, and transposed versions shifted by a semitone and a whole tone were created. For each version, piano-roll images were generated and segmented into multiple two-measure vectors, which were used as inputs to the visible layer. Corresponding hidden-unit activations were sampled and analyzed.
Figure 11 shows the results of dimensionality reduction of the hidden activations using t-SNE [Van der Maaten and Hinton, 2008]. In the case of semitone transposition, the hidden representations of the original and transposed versions were distributed at relatively distant locations in the low-dimensional space (Fig. 11(a)). This indicates that the hidden states changed substantially after transposition, suggesting that the RBM treated the transposed data as distinct inputs. In contrast, when the pieces were transposed by a whole tone, the hidden representations of the original and transposed versions were located in closer proximity (Fig. 11(b)).
@@
To further interpret this behavior, the number of shared scale tones between the original and transposed keys was examined. For standard seven-note scales (e.g., major and natural minor), a semitone transposition shares only two scale tones with the original key, whereas a whole-tone transposition shares five tones. The results suggest a tendency for hidden-state vectors to be located closer together when the transposed and original inputs share a larger number of scale tones.
@@
These observations indicate that the RBM primarily encodes absolute pitch information rather than relative pitch relationships when evaluating similarity. During training, the dataset was augmented by including transpositions of the original pieces up to a major sixth upward and a perfect fourth downward, analogous to data augmentation in image processing. Despite this, the trained RBM remained sensitive to pitch translations and did not exhibit transposition invariance. This behavior is consistent with previous reports that RBMs are vulnerable to translations in the input space [Lee et al., 2009] and was also observed in the present piano-roll experiments.
Refer to caption(b)
@@
Appendix B GPGPU-Accelerated Training of the RBM
We developed a lightweight Python library designed to facilitate systematic experimentation with restricted Boltzmann machines (RBMs). The library is publicly available [Kobayashi and Watanabe, 2025] and was implemented with a focus on portability and transparency. It performs numerical computations using NumPy when executed on a CPU and automatically switches to CuPy when a compatible GPGPU is detected, thereby enabling hardware-accelerated computation without requiring changes to user code.
To characterize the practical performance of this implementation, we measured the training time of RBMs on the MNIST dataset using System C, a supercomputer at the Institute for Solid State Physics, The University of Tokyo. Computation times were evaluated both in CPU-only mode and in GPGPU-accelerated mode. The CPU was an AMD EPYC 7763 (2.45 GHz, 64 cores) with 256 GB of main memory, and the GPGPU configuration consisted of four NVIDIA A100 GPUs (40 GB memory each, total 160 GB). Training times were measured for varying numbers of hidden units, and the results are summarized in Fig. 12.
The results show that the library successfully utilizes GPGPU acceleration and achieves substantial reductions in training time compared to CPU-only execution. These measurements confirm that the implementation supports scalable, hardware-accelerated experimentation with RBMs and is suitable for large-scale model exploration.